OpenStudy (pphalke):

Nico is saving money for his college education. He invests some money at 5% and $1900 less than the amount at 3%. The investments produced a total of $143 interest in 1 yr. How much did he invest at each rate?? @ILovePuppiesLol @adrynicoleb

ILovePuppiesLol (ilovepuppieslol):

Hello, we're sorry for the inconvenience but us helpers only give help as volunteers, we don't get anything in exchange. That is why there are not a lot of us around... Please be patient and be as clear as you can with your questions :)

ILovePuppiesLol (ilovepuppieslol):

u can tag others if u want!!

ILovePuppiesLol (ilovepuppieslol):

I'm busy atm :(

OpenStudy (pphalke):

ok

OpenStudy (welshfella):

let hom invest $x at 5% so he invests x - 1900 at 3 %

OpenStudy (welshfella):

after 1 year he gets $0.05x at 5% adn $ (x - 1900)0.03 at 3% these 2 add up to $143 so we have the equation 0.05x + 0.03(x - 1900) = 143 can you solve this for x?

OpenStudy (welshfella):

first distribute the 0.03 over the parentheses; 0.05x + 0.03x - 0.03*1900 = 143 0.08x - 57 = 143 can you continue and find the value of x?

OpenStudy (welshfella):

add 57 to both sides of the equation 0.08x - 57 = 143 +57 +57 0.08x = 200 now divide right side by 0.08 and you have the values of x , the amount invested at 5%

OpenStudy (pphalke):

i got 2500 out of that

OpenStudy (welshfella):

thats correct so all you need do to find amoubt invested at 3* is subtract 1900 from 2500

OpenStudy (pphalke):

600 invested at 3%

OpenStudy (welshfella):

yep

OpenStudy (welshfella):

ok? any questions?

OpenStudy (pphalke):

i have more ill post them